OpenStudy (anonymous):

Simplify the expression 36^-1/2. answers: 1/18 1/6 -6 -18..

OpenStudy (anonymous):

As I said: \[a^{-b} = \frac{1}{a^b}\] Here, a = 36 and b = 1/2 Put the values and solve then..

OpenStudy (anonymous):

\[36^{-1/2}=1/36^{1/2}=1/6\]

OpenStudy (anonymous):

thanks :)